OpenStudy (anonymous):

2n+60=90 solve for n...is n 15?

jimthompson5910 (jim_thompson5910):

n = 15 is correct

jimthompson5910 (jim_thompson5910):

Check: 2n + 60 = 90 2*15 + 60 = 90 ... replace all copies of n with 15 30 + 60 = 90 90 = 90 ... answer is confirmed

OpenStudy (anonymous):

N= 15.

OpenStudy (anonymous):

90 is wrong for n